OpenStudy (anonymous):

how in the world do I figure this out, please explain in detail. The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? 56 41 69 55

OpenStudy (jdoe0001):

well a+ b = 110 a * b = 28 so, is a system of equations use substitution say, on the 2nd one, if you solve for "a", what would that be?

OpenStudy (marissalovescats):

Their difference is 28 not their product @jdoe0001

OpenStudy (campbell_st):

well actually the sum is a + b = 110 a - b = 28 add the 2 equations together which will allow you to eliminate b 2a = 138 solve for a... this will be the larger of the 2 numbers

OpenStudy (jdoe0001):

ohh darn, ok

OpenStudy (jdoe0001):

misread there

OpenStudy (primeralph):

Can I use Number Theory? Please? 

OpenStudy (jdoe0001):